OpenStudy (anonymous):

Any help is appreciated! Solve the equation for x. Identify any extraneous solutions. √x=-3 I know the answer is (+9) but is (-9) an extraneous solution? I'm not sure what an extraneous solution is. Thanks!

OpenStudy (anonymous):

an extraneous solution is basically a value that seems to work when solving an equation, when in reality it doesn't work

OpenStudy (anonymous):

for this problem, you can't have a negative value under a square root sign, so -9 should not be a solution

OpenStudy (anonymous):

So it is NOT an extraneous solution? @11calcBC

OpenStudy (anonymous):

if my memory is accurate, yes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

somebody confirm this...

OpenStudy (anonymous):

Can you confirm this @tkhunny ?

OpenStudy (tkhunny):

\(\sqrt{x} = -3\) You do NOT know the answer is x = +9. That is incorrect. The answer is also not x = -9. This equation has no solution in the Real Numbers. The Domain if \(f(x) = \sqrt{x}\) is \(x \ge 0\). This eliminates \(x = -9\) The Range if \(f(x) = \sqrt{x}\) is \(f(x) \ge 0\). This eliminates \(x = 9\)

OpenStudy (tkhunny):

\(\sqrt{9} = 3 \ne -3\)

OpenStudy (anonymous):

Thanks again @11calcBC !!!

OpenStudy (anonymous):

no problem :D

OpenStudy (tkhunny):

What was your final answer?